OpenStudy (anonymous):

If f(x) varies directly with x2, and f(x) = 72 when x = 6, find the value of f(3). A. 18 B. 54 C. 36 D. 108

OpenStudy (anonymous):

f(x) = Kx^2 72 = K6^2 K = 72/6^2 = 2 f(x) = 2x^2 f(3) = 2*3^2 f(3) = 18

OpenStudy (anonymous):

So its 18?

OpenStudy (anonymous):

Yes it's 18